OpenStudy (anonymous):

Suppose a 10 ft/ long ladder is placed on the ground agaisnt a 5ft wall, so as to exactly reach the top of the wall. What angle does the ladder make with the ground ? I'm supposed to get 30 degrees. I know there's the 90, 60, 30 triangle yet how am I supposed to know if I pick 60 or 30 ?

OpenStudy (anonymous):

its 30 degrees sin^{-1} (5/10)

OpenStudy (anonymous):

@elsanna since I can't use a calculator on the test I would just need to know that a triangle is always 30 degrees at the base ?

OpenStudy (anonymous):

your wall height is 5 ft and ladder is 10ft. if you draw a figure it looks like this

OpenStudy (anonymous):

Nevermind I see why 5/10 is 1/2 and sin^-1 =1/2 is the same as sin=1/2 which is 30 degrees, and since the ladder was set as opp/adj its sin, right ?

OpenStudy (anonymous):

I mean opp/ hyp

OpenStudy (anonymous):

Thank you so much :)